HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ATTOSECOND PULSES FOR STUDYING ULTRAFAST ELECTRON DYNAMICS

\underline{A. L’HUILLIER} , Physics Department, Lund University

Extreme Ultraviolet light sources based on high-order harmonic generation in gases are now used in many areas of science. The radiation consists of a train of extremely short light bursts, in the 100-attosecond range, allowing for outstanding temporal resolution.

Attosecond pulses have enabled the study of photoionization of atoms and molecules in a completely new way. Through the availability of a synchronized probe field and interferometric measurements, it has become possible to measure the incredibly small-time delay in photoionization, a concept introduced by E. Wigner in 1955, and the phase change of an electron wavepacket across a Fano resonance. Recently, the use of a bicolor probe field has allowed us to perform the quantum state tomography of a photoelectron.

p-number: p246  Submitted on Mon, 26 Jun 2023 09:12:22 +0200